# Digital Twin-Enabled Real-Time Control in Robotic Additive Manufacturing via Soft Actor-Critic Reinforcement Learning

### v1

@@ Abstract
Smart manufacturing systems increasingly rely on adaptive control mechanisms to optimize complex processes. This research presents a novel approach integrating Soft Actor-Critic (SAC) reinforcement learning with digital twin technology to enable real-time process control in robotic additive manufacturing. We demonstrate our methodology using a Viper X300s robot arm, implementing two distinct control scenarios: static target acquisition and dynamic trajectory following. The system architecture combines Unity’s simulation environment with ROS2 for seamless digital twin synchronization, while leveraging transfer learning to efficiently adapt trained models across tasks. Our hierarchical reward structure addresses common reinforcement learning challenges including local minima avoidance, convergence acceleration, and training stability. Experimental results show rapid policy convergence and robust task execution in both simulated and physical environments, with performance metrics including cumulative reward, value prediction accuracy, policy loss, and discrete entropy coefficient demonstrating the effectiveness of our approach. This work advances the integration of reinforcement learning with digital twins for industrial robotics applications, providing a framework for enhanced adaptive real-time control for smart additive manufacturing process.
@@ I Introduction
The evolution of smart manufacturing demands increasingly sophisticated control systems capable of adapting to dynamic production environments. Reinforcement learning (RL) has emerged as a promising approach for developing such adaptive control systems, offering advantages over traditional control methods through its ability to optimize behavior through environmental interaction [1]. Unlike supervised learning approaches that rely on labeled datasets, RL enables autonomous exploration and strategy refinement, making it particularly valuable for robotics applications where continuous adaptation is essential [2, 3].
@@
However, implementing RL in manufacturing environments presents several significant challenges. These include sample inefficiency during training, computational intensity, and potential instability in learned policies [4]. Additionally, the direct application of RL algorithms in physical manufacturing systems risks equipment damage and production disruption during the learning phase [5, 6]. Recent applications have demonstrated RL’s potential across various manufacturing domains. Wang et al. [7] successfully applied RL to optimize CNC machining parameters in real-time, achieving improvements in surface quality while reducing tool wear. In the automotive sector, Loffredo et al. [8] implemented RL-based energy-efficient control systems for multi-stage production lines, effectively balancing energy consumption with throughput requirements. Moreover, RL methods are particularly effective in improving the robustness of microgrids, such as DC microgrids integrating piezoelectric energy harvesting, which shows substantial improvement under failure scenarios [9].
In additive manufacturing, RL has shown promise for process optimization under varied manufacturing environments. Li et al. [10] developed an RL-based predictive model for surface roughness in fused filament fabrication, while Ge et al. [11] utilized RL to optimize path planning for thin-walled structure printing. These studies highlight RL’s capacity to enhance manufacturing precision and efficiency, though challenges persist in areas such as real-time adaptation and multi-objective optimization.
Digital twin technology offers a potential solution to these implementation challenges by providing a safe, virtual robotic additive manufacturing environment for RL training and real-time communication with real robot path planning for 3D printing process. Digital twins enable real-time mirroring of physical systems, facilitating simulation-based optimization and monitoring [12]. When combined with RL, digital twins create a protected testing environment that minimizes physical risks while allowing for rapid iteration and validation [13].
Furthermore, the digital twin synchronization setup enables robust RL model training in a virtual environment, where trained models can then be validated in real-world settings. This integration provides a high-fidelity testing environment for robotic control algorithms, improving safety, reliability, and efficiency in robotics research [14, 15, 16, 17].
I-A Unity’s Role in Simulation and Real-Time Control
Unity has emerged as a valuable simulation platform across fields such as robotics, autonomous navigation, temperature control, and underwater vehicle testing. Its integration with Robot Operating Systems (ROS and ROS2) enables real-time control and synchronization between virtual simulations and physical robots, making it a preferred choice for simulating RL-based control systems. In collaborative robotics and motion planning, Unity allows for interactive path editing and real-time trajectory generation, which is essential for optimizing robot control strategies in hazardous or dynamic environments [18]. The applications with Unity have underscored Unity’s value for autonomous systems development and real-time control, where high-fidelity virtual environments allow for extensive testing and refinement of RL algorithms without the constraints of real-world testing [19, 20].
The Temporal Gauss-Seidel (TGS) solver in Unity enhances the realism and stability of physics simulations by iteratively resolving constraints, stabilizing interactions, and reducing jitter, particularly in real-time applications. TGS’s efficiency in managing complex collisions and joint interactions is crucial for scenarios requiring precision, such as robotics, vehicle dynamics, and immersive virtual environments. For example, TGS supports applications like real-time robot control and vehicle navigation by stabilizing simulations to reflect realistic physical behavior [21]. This solver provides computational efficiency, making Unity a top choice for interactive environments where both realism and real-time performance are essential. Other approaches, such as physics-based deep learning frameworks, offer alternative methods that may provide more generalized solutions for broader applications, but Unity’s TGS solver remains a practical choice for scenarios requiring immediate, reliable responses [22, 23].
I-B Challenges in Robot Arm Control and Path Optimization
@@
Robot arm control and path optimization present several challenges, especially in dynamic environments where the task conditions are constantly changing [24]. Traditional control methods often struggle to adapt in real-time, relying on predefined models and static decision-making processes. Path optimization is particularly challenging because robots must calculate the most efficient trajectory while avoiding obstacles, complying with kinematic constraints, and reacting to real-time environmental feedback [25]. These factors make path planning in robotics a complex and time-consuming task.
Reinforcement learning offers a potential solution to these challenges, as it enables robots to learn optimal strategies for controlling motion and planning paths. By continuously interacting with the environment and receiving feedback, RL agents can improve their decision-making over time, leading to better performance in dynamic settings. However, RL-based systems must overcome issues such as slow convergence, local minima, and instability in learning, which can limit their effectiveness. This research aims to address these challenges by leveraging the Soft Actor-Critic (SAC) RL algorithm and Hierarchical Reward Structure (HRS) to improve the training efficiency, stability, and adaptability of robotic systems, particularly in path optimization tasks.
The primary objective of this research is to explore the integration of reinforcement learning (specifically the Soft Actor-Critic algorithm) with digital twin technology to enhance robot arm control and optimize path planning. The key goals of the study are outlined as follows:
@@
By combining Unity’s simulation environment, RL training, ROS2, and real robot arm communication, this research aims to advance robotic adaptability, task stability, and learning efficiency, contributing to the broader fields of autonomous robotics applications and smart additive manufacturing process.
@@ II-A Experimental Setup and Digital Twin Synchronization
II-A1 Viper X300s Robot Arm
Our experimental setup utilizes the Viper X300s, a Six Degree of Freedom (DOF) robot arm designed by Trossen Robotics for research applications. With a span of 1500mm and a payload capacity of 750g, the Viper X300s offers researchers a compact yet powerful platform for experimenting with robotic functionalities. Through ROS inter-bridging, researchers can simulate tasks in Unity’s virtual environment and then validate them on the Viper X300s, creating a seamless feedback loop between virtual testing and physical deployment.
The Viper X300s, designed with biologically inspired adaptability, can operate effectively in unstructured environments, making it well-suited for applications like exploration, emergency response, and even hazardous material handling. It features low-cost computer vision capabilities combined with neural networks for object tracking and classification, which broadens its adaptability across various automation tasks, from precision assembly to search and rescue operations [26, 27].
II-A2 Digital Twin Unity Environment Setup
The Unified Robot Description Format (URDF) file for the Viper X300s robot arm was imported into Unity’s virtual environment, which included a simulated desk setup, enabling accurate simulations of the robot’s kinematics and movement dynamics. This allows researchers to manipulate joint positions, adjust trajectories, and experiment with spatial orientation, providing a practical way to simulate robotic behavior before real-world application. By integrating polynomial algorithms, the simulation can achieve high precision in task execution, helping to ensure accurate and repeatable operations [28]. Simulating tasks like pick-and-place within Unity reduces the wear on physical hardware, allowing for extensive testing without physical risk. Unity also supports collaborative, remote-access features, making the virtual environment an accessible and versatile tool for robotics research and training in a controlled, risk-free setting [29].
However, there is a lack of ROS2 package for subscribing to Unity Publisher for the Viper X300s robot arm, hence a ROS2 subscriber package was developed. Establishing digital twin connection as given in 1, where virtual viper x300s robot arm in Unity continuously sends data of the robot arms joints to a local TCP-server and the developed ROS2 package subscribes to these data and move the robot arm continuously.
II-A3 ROS-TCP Connector for Communication and Control
The Unity ROS-TCP-Connector extension enables seamless publishing of ROS2 messages to a local cloud server, allowing bidirectional communication between Unity and ROS-based systems. These messages are subscribed by ROS2 for processing on the physical Viper X300s robot arm, establishing real-time digital twin synchronization. This synchronization has a latency of only about 20 milliseconds, ensuring near-instantaneous reflection of virtual movements on the physical counterpart, which allows for testing and validation of robotic tasks in a simulated environment before deployment [30, 31, 32].
@@ II-B1 Unity ML-Agents Framework for Reinforcement Learning
Unity’s ML-Agents extension enables RL-based agent training in simulated environments, including robotics, game AI, and autonomous systems. We implemented the RL framework as shown in Figure 2. With support for RL algorithms such as Proximal Policy Optimization (PPO) and Soft Actor-Critic (SAC), the ML-Agents toolkit facilitates complex simulations where agents learn through continuous interaction with their environment. ML-Agents allows Unity to act as a simulation environment while handling training processes in Python, creating an efficient workflow for testing and optimizing RL models in robotics [33, 34].
@@
Our established algorithms environment was run in Unity that consisted of different objects and a virtual robot arm as illustrated in Figure 2. The algorithms were trained through Python Trainer. Initially, the states of the environment was fed to the RL agent so that it could gain an understanding of the working environment. After getting the state, the agent took a set of action policies to asses the change in state and reward. These changes were then further fed into the agent to store the set of action policies in replay buffer. Following this implementation, the agent improved its set of action policies and then evaluated them. After which the environment communicated the state and reward back to the agent. This iteration was done throughout the training, and after the Case 1 training was completed, the trained model was transferred to the case 2 training which also utilized the trained models’ weights to adjust and improve its learning. This training process has been shown in Figure 3.
@@
In our research, we utilized the SAC algorithm due to its faster convergence speed compared to PPO, making it suitable for tasks that require rapid learning and adaptation. The SAC algorithm balances exploration and exploitation by storing agent actions and corresponding rewards in a replay buffer, enabling the agent to iteratively improve its policy for maximum cumulative reward. Hyperparameter tuning was used to optimize performance by adjusting parameters such as learning rate and exploration rate to prevent local minima during training [35, 36]. The Berkeley AI Research (BAIR) blog [37] developed SAC’s advantages in control tasks, which guided our decision to select this algorithm [38].
@@ II-B2 Action Space for Robot Arm Control
The choice of dynamic model and controller greatly influences the efficiency of training. Different dynamic models, including kinematic and dynamic models, can affect both the accuracy and speed of RL tasks in robotic navigation [39]. Moreover, the use of sliding mode control (SMC) in trajectory tracking, particularly in challenging environments with sharp turns, can benefit from trajectory smoothing techniques that improve RL performance [40]. Additionally, RL can be applied in microgrids for voltage and frequency regulation, demonstrating its ability to outperform conventional controllers and improve system stability in islanded conditions [41].
@@ II-B3 Hierarchical Reward Structure Policy
A hierarchical reward structure policy was implemented to improve learning efficiency.
The tables LABEL:tab:Case1, LABEL:tab:Case2 and LABEL:tab:Case3 below outlines the reward framework designed for two cases in the reinforcement learning task, highlighting the primary goals, sub-goals, reward values, and the utilization of transfer learning to enhance model adaptation. The listed framework strategically combines positive reinforcement for achieving objectives and maintaining stability with penalties for undesirable actions, while leveraging transfer learning from simpler tasks to more complex ones.
This strategy decomposes the task into sub-goals, each with its own reward structure, allowing the agent to achieve incremental successes and maintain consistent learning progress. By incorporating intermediate rewards for sub-goals, the agent learns high-level strategies and low-level actions more effectively, improving training scalability and stability [42, 43]. Sub-goals in RL tasks enhance convergence speed and task accuracy [44], while structured reward policies in multimodal systems manage data fusion complexity, boosting efficiency and accuracy [45, 46]. Hierarchical models in VR training systems for manufacturing also optimize learning outcomes by adapting to novice learners’ pace [47].
To further accelerate learning, we used transfer learning based on the hierarchical reward structure, where the trained model from Case 1 was fine-tuned for Case 2, leveraging pre-learned features to reduce training time and computational costs. Transfer learning enhances performance and efficiency, particularly in scenarios where the new task shares similarities with a previously trained model.
@@ III Results and Discussion
The experimental results demonstrate the effectiveness of our SAC-based reinforcement learning approach across multiple performance metrics, as illustrated in Figure 4. The analysis encompasses four key dimensions: cumulative reward, episode length, policy loss, and value prediction accuracy, tracked over 200,000 training steps.
@@ III-A1 Cumulative Reward Progression
Figure 4 reveals distinct learning patterns across three experimental cases. In Case 1 (static target-reaching), the agent initially encountered a significant local minimum, evidenced by the sharp dip in reward percentage around the 25,000-step mark. However, the implementation of our hierarchical reward structure enabled recovery, leading to stable convergence at approximately step of 60,000. Case 3 without the hierarchical transfer training showed a gradual reward increase but with slow convergence until the step of 150,000.
Case 2, leveraging transfer learning from Case 1, demonstrated remarkably faster convergence despite handling the more complex task of following a moving target. This acceleration validates our transfer learning approach which showed clear improvements. Case 2 also achieved gradual increase of reward value, at the same time with fast speed of convergence at a step of 50,000, suggesting superior policy performance compared to Cases 1 and 3. Additionally, Case 2 reached a higher cumulative reward more quickly and maintained it consistently, achieving the task’s goals more effectively than Case 3.
@@ III-A3 Policy Evolution and Entropy Analysis
Policy loss trends in Figure 4 reveal the adaptation characteristics of the SAC algorithm. It illustrates the tendency of the agent to change its setup actions per iteration. Case 1’s policy loss stabilized at -200%, while Case 2 showed a gradual decline to -1200%, reflecting increasingly deterministic policy selection. Figure 4 further revealed that Cases 2 and 3 experienced greater policy adjustments, as seen by their large negative values, while Case 1 remains relatively stable. Although both cases show substantial negative policy loss, Case 2’s loss stabilizes faster, implying that it found an effective policy earlier and with less overcorrection than Case 3. The discrete entropy coefficient from Figure 4 further supports this observation, demonstrating the agent’s transition from exploration to exploitation, particularly evident in Case 2’s rapid entropy reduction after 50,000 steps. Case 3 took the longest to stabilize, taking approximately 125,000 steps. Case 2 initially encouraged exploration with a high entropy coefficient but reduced it faster than Case 3, signaling a more efficient transition from exploration to exploitation.
@@ III-B Real-world Validation Through Digital Twin Synchronization for Robotic AM Line Scanning
The success of the hierarchical training structure is largely attributed to several key factors. First, the use of high-fidelity physics simulation, facilitated by Unity’s TGS solver, ensures accurate and realistic modeling of physical interactions. Second, the implementation of a robust reward structure effectively incorporates practical constraints, guiding the agent toward achieving desired behaviors while penalizing undesirable actions. Lastly, effective domain randomization during training enhances the model’s adaptability and generalization, enabling it to perform reliably under varying real-world conditions.
The results demonstrate enhanced stability in transfer learning, with a notably smooth transition between Cases 1 and 2. This stability indicates that the SAC-based approach can maintain consistent performance when transferring learned policies across tasks, which is essential for applications requiring adaptive control.
Furthermore, Case 2 achieved higher training efficiency and improved overall performance compared to Case 3, emphasizing the effectiveness of integrating transfer learning within this framework. The reduced training time and increased reward attainment in Case 2 suggest that transfer learning not only accelerates convergence but also helps the model adapt more effectively to task variations. This efficiency is especially valuable in robotic systems, where rapid learning and adaptation are critical for handling dynamic and evolving tasks. These improvements can be attributed to our hierarchical reward structure and the effective integration of transfer learning principles.
@@ IV Conclusion
Our research successfully demonstrates the integration of RL with digital twin technology to enhance robotic control within manufacturing applications. By utilizing the SAC algorithm and the Viper X300s platform, we tackle several key challenges in smart manufacturing, ultimately developing adaptive and efficient control systems. The findings from this work highlight the transformative potential of RL-driven digital twins, showcasing their ability to enable robust, real-time synchronization between virtual simulations and physical additive manufacturing process. This breakthrough opens new avenues for optimizing automation processes and bridging the gap between simulation and real-world applications.
IV-A Contributions and Performance Analysis
The primary contributions of this study include the development of a high-performance digital twin synchronization framework that achieves approximately 20ms latency, the implementation of a hierarchical reward structure that significantly enhances learning efficiency and policy stability, and the successful application of transfer learning techniques across related tasks to drastically reduce training time. In terms of performance, the SAC algorithm demonstrated its capability to achieve consistent convergence in static tasks within just 60,000 steps, while transfer learning enabled rapid adaptation to more complex dynamic tasks with minimal additional training. Notably, when transferring the learned policies to physical hardware, performance degradation was limited to less than 5%. These results validate the effectiveness of using digital twins for rapid prototyping, testing control strategies, and exploring edge cases—all while minimizing the need for extensive real-world testing, and consequently reducing system downtime. These achievements underscore the potential of RL and digital twin integration in advancing automation within manufacturing settings.
IV-B Future Work
Despite these promising results, scalability, environmental variability, and real-time adaptation remain challenges. Future work will focus on extending the framework to multi-robot coordination, investigating hybrid learning approaches such as combining SAC with imitation learning or model-based RL, and enhancing digital twin fidelity to narrow the reality gap. Additionally, research will explore seamless integration with manufacturing execution systems and enterprise resource planning platforms. These advancements aim to expand the applicability of RL and digital twin technologies in adaptive, high-precision, and flexible manufacturing systems.
The findings of this study demonstrate that reinforcement learning, when effectively integrated with digital twin technology, provides the adaptability and reliability needed for next-generation manufacturing systems. Success in both static and dynamic tasks, combined with efficient transfer learning capabilities, underscores the potential of this approach. These results pave the way for more autonomous and adaptable manufacturing processes, establishing a solid foundation for future innovations in smart manufacturing

### v2

# Digital Twin Synchronization: Bridging the Sim-RL Agent to a Real-Time Robotic Additive Manufacturing Control

@@ Abstract
With the rapid development of deep reinforcement learning technology, it gradually demonstrates excellent potential and is becoming the most promising solution in the robotics. However, in the smart manufacturing domain, there is still not too much research involved in dynamic adaptive control mechanisms optimizing complex processes. This research advances the integration of Soft Actor-Critic (SAC) with digital twins for industrial robotics applications, providing a framework for enhanced adaptive real-time control for smart additive manufacturing processing. The system architecture combines Unity’s simulation environment with ROS2 for seamless digital twin synchronization, while leveraging transfer learning to efficiently adapt trained models across tasks. We demonstrate our methodology using a Viper X300s robot arm with the proposed hierarchical reward structure to address the common reinforcement learning challenges in two distinct control scenarios. The results show rapid policy convergence and robust task execution in both simulated and physical environments demonstrating the effectiveness of our approach.
@@ I Introduction
As a significant aspect of artificial intelligence, reinforcement learning (RL) provides an approach to helping robots develop self-learning capability. RL can increase the adaptability of robotics systems, which is an important feature in order to deal with a complex and dynamic environment [1, 2]. However, there are many challenges to implementing RL in the robotic domain due to hardware limitations (such as mechanical system design and computer capabilities) and software restrictions (like algorithm efficiency). Reinforcement learning (RL) has emerged as a promising approach for developing such adaptive control systems, offering advantages over traditional control methods through its ability to optimize behavior through environmental interaction [3]. Unlike supervised learning approaches that rely on labeled datasets, RL enables autonomous exploration and strategy refinement, making it particularly valuable for robotics applications where continuous adaptation is essential.
@@
From the manufacturing perspective, the challenges include sample inefficiency during training, computational intensity, and potential instability in learned policies [4]. Additionally, the direct application of RL algorithms in physical manufacturing systems risks equipment damage and production disruption during the learning phase. Recent applications have demonstrated RL’s potential across various manufacturing domains. Wang et al. [5] successfully applied RL to optimize CNC machining parameters in real-time, achieving improvements in surface quality while reducing tool wear. In the automotive sector, Loffredo et al. [6] implemented RL-based energy-efficient control systems for multi-stage production lines, effectively balancing energy consumption with throughput requirements.
@@
However, robotic arm control and path optimization in industrial manufacturing rely on predefined models and static decision-making processes. Especially in path planning, robots must calculate the most efficient trajectory while avoiding obstacles, complying with kinematic constraints, and reacting to real-time environmental feedback, which makes path optimization a complex and time-consuming task.
To address this gap, we propose a digital twin synchronization method that transforms the well-trained simulation RL (sim-RL) agent through Unity into a real robot arm in real time within the Unity and ROS2 framework. Moreover, to overcome RL challenges such as slow convergence, local minima, and instability in the training process, we introduce the Hierarchical Reward Structure (HRS) integrating with the Soft Actor-Critic (SAC) [7] method to achieve sample efficiency. The main contributions of this research is listed as below.
@@ II-A Experimental Setup and Digital Twin Synchronization
Our experimental setup utilizes the Viper X300s, a Six Degree of Freedom (DOF) robot arm designed by Trossen Robotics for research applications. With a span of 1500mm and a payload capacity of 750g, the Viper X300s offers researchers a compact yet powerful platform for experimenting with robotic functionalities. Through ROS inter-bridging, researchers can simulate tasks in Unity’s virtual environment and then validate them on the Viper X300s, creating a seamless feedback loop between virtual testing and physical deployment. We built a ROS2 package to receive data from the virtual model of the Viper X300s robot arm in Unity through a local TCP server. Fig. 1 illustrates the establishing digital twin connection. The Unity ROS-TCP-Connector extension enables seamless publishing of ROS2 messages to a local cloud server, allowing bidirectional communication between Unity and ROS-based systems. These messages are subscribed by ROS2 for processing on the physical Viper X300s robot arm, establishing real-time digital twin synchronization. This synchronization has a latency of only about 20 milliseconds, ensuring near-instantaneous reflection of virtual movements on the physical counterpart, which allows for testing and validation of robotic tasks in a simulated environment before deployment [8].
@@ II-B1 Unity ML-Agents Framework for Reinforcement Learning
Our established algorithms environment was run in Unity [9] consisting of different objects and a virtual robot arm as illustrated in Fig. 2. The algorithms were trained through Python Trainer. Initially, the states of the environment were fed to the RL agent so that it could gain an understanding of the working environment. After getting the state, the agent took a set of action policies to assess the change in state and the reward. These changes were further fed into the agent to store the action policies in the replay buffer. Following this implementation, the agent improved its set of action policies and then evaluated them. After that, the environment communicated the state and reward back to the agent. This iteration was done throughout the training, and after the Case 1 training was completed, the trained model was transferred to the Case 2 training, which also utilized the trained models’ weights to adjust and improve its learning. The training process has been shown in Fig. 3.
@@ II-B3 Hierarchical Reward Structure Policy
We design a hierarchical reward structure policy to improve learning efficiency. The tables LABEL:tab:Case1, LABEL:tab:Case2 and LABEL:tab:Case3 outline the reward framework designed for two cases in the reinforcement learning task, highlighting the primary goals, sub-goals, reward values, and the utilization of transfer learning to enhance model adaptation. The listed framework strategically combines positive reinforcement to achieve objectives and maintain stability with penalties for undesirable actions, while leveraging transfer learning from simpler tasks to more complex ones. This strategy decomposes the task into sub-goals, each with its own reward structure, allowing the agent to achieve incremental successes and maintain consistent learning progress. By incorporating intermediate rewards for sub-goals, the agent learns high-level strategies and low-level actions more effectively, improving training scalability and stability. To further accelerate learning, we used transfer learning based on the hierarchical reward structure, where the trained model from Case 1 was fine-tuned for Case 2, leveraging pre-learned features to reduce training time and computational costs. Transfer learning enhances performance and efficiency, particularly in scenarios where the new task shares similarities with a previously trained model.
@@ III Results and Discussion
The analysis encompasses four key dimensions: cumulative reward, episode length, policy loss, and value prediction accuracy, tracked over 200,000 training steps.
@@ III-A1 Cumulative Reward Progression
Figure 4 reveals distinct learning patterns across three experimental cases. In Case 1 (static target-reaching), the agent initially encountered a significant local minimum, evidenced by the sharp dip in reward percentage around the 25,000-step mark. However, the implementation of our hierarchical reward structure enabled recovery, leading to stable convergence at approximately step of 60,000. Case 3 without the hierarchical transfer training showed a gradual reward increase but with slow convergence until the step of 150,000. Case 2, leveraging transfer learning from Case 1, demonstrated remarkably faster convergence despite handling the more complex task of following a moving target. Additionally, Case 2 reached a higher cumulative reward more quickly and maintained it consistently, achieving the task’s goals more effectively than others.
@@ III-A3 Policy Evolution and Entropy Analysis
Policy loss trends in Figure 4 reveal the adaptation characteristics of the SAC algorithm. It illustrates the tendency of the agent to change its setup actions per iteration. Case 1’s policy loss stabilized at -200%, while Case 2 showed a gradual decline to -1200%, reflecting increasingly deterministic policy selection. Figure 4 further revealed that Cases 2 and 3 experienced greater policy adjustments, as seen by their large negative values, while Case 1 remains relatively stable. Although both cases show substantial negative policy loss, Case 2’s loss stabilizes faster, implying that it found an effective policy earlier and with less overcorrection than Case 3. The discrete entropy coefficient from Figure 4 further supports this observation, demonstrating the agent’s transition from exploration to exploitation, particularly evident in Case 2’s rapid entropy reduction after 50,000 steps.
@@ IV Conclusion
This research demonstrates the integration of RL with digital twin technology to enhance robotic control within manufacturing applications. By utilizing the SAC algorithm and the Viper X300s platform, we tackle several key challenges in smart manufacturing, ultimately developing adaptive and efficient control systems. The findings from this work highlight the transformative potential of RL-driven digital twins, showcasing their ability to enable robust, real-time synchronization between virtual simulations and physical additive manufacturing process. This breakthrough opens new avenues for optimizing automation processes and bridging the gap between simulation and real-world applications.